OpenStudy (zeesbrat3):

Which of the following statements is/are true? a.) If f '(c) = 0, then f has a local maximum or minimum at x = c. b.) If f is continuous on [a, b] and differentiable on (a, b) and f '(x) = 0 on (a, b), then f is constant on [a, b]. c.) The Mean Value Theorem can be applied to f(x) = 1/x2 on the interval [−1, 1].

OpenStudy (zeesbrat3):

@ganeshie8 @mathstudent55 @kropot72 @Kash_TheSmartGuy

OpenStudy (zeesbrat3):

Please help

OpenStudy (michele_laino):

hint: for last statement, we have that x=0 belongs to interval [-1,1], and at x=0, the functio f(x)= 1/x^2 is not continuous

OpenStudy (zeesbrat3):

Because it wouldnt exist?

OpenStudy (michele_laino):

since we can not divide by zero

OpenStudy (zeesbrat3):

Sorry, I thought of limits when I said it doesn't exist since like you said, you can't divide by 0

OpenStudy (zeesbrat3):

I think A i incorrect as well

OpenStudy (michele_laino):

yes! I think so, since we can have this case: 

OpenStudy (zeesbrat3):

I did it algebraically, but got the same thing. I figured if you had f(x) = x^3 then you wouldn't get a max or min anyways, just a hip

OpenStudy (zeesbrat3):

So, since A and C are incorrect, the only valid answer would be B, yes?

OpenStudy (michele_laino):

another counterexample is given by this function: f(x)= x^3 we have: f '(x) = 3x^2 and f '(0)==, nevertheless x=0 is not a point of local maximum or minimum